A support cable for a 64 foot tower forms a 70° angle with the ground. What is the length of the cable to the nearest foot? Thankx
Answer by arunpaul(104)

We know Sin@= hypertnews/perpendicular = h/l
so sin70° = x/64
let x = h= length of the support cable
0.9396 = (x/64)
or x = .9396*64
or x = 60.1403 ft (Ans)